I have tried flipping the problem to make it in positive exponent form, that has not helped me. I tried reducing terms first, not sure if I should do this. The problem solver on the website does not give the correct answer, as I am able to check it because it is in the back of the book. I cannot do the correct steps to find
(125a^2b^-3/5a^4b^-2)^-5 = a^10b^5/5^10

Answer by Alan3354(30993)

(125a^2b^-3/5a^4b^-2)^-5 = a^10b^5/5^10
=============
(125a^2b^-3/5a^4b^-2)^-5
= (25a^2b^-3/a^4b^-2)^-5
= (25b^-3/a^2b^-2)^-5
= (25b^-1/a^2)^-5
= (25/a^2b)^-5
= (a^2b/25)^5
= (a^2b/5^2)^5
= a^10b^5/5^10